How much 1% boric acid solution and 5% boric acid solution are needed to make 30 mL of 3% boric acid solution?

Answer by mananth(16949)

How much 1% boric acid solution and 5% boric acid solution are needed to make 30 mL of 3% boric acid solution?
let 1% boric acid solution required be x ml
The concentration in the mixture will be 1/100* x = 0.01x
The balance 30-x ml will be 5% solution
The concentration of this in the mixture will be 0.05*(30-x)
The concentration of the total mixture will be 0.03 * 30
0.01x+0.05(30-x)= 0.03*30
0.01x+ 1.5 - 0.05x = 0.9
-0.04x= -0.6
x=0.6/0.04
x=15 ml 1 % solution
balance 15 ml will be 3% solution